# SafeProtein: Red-Teaming Framework and Benchmark for Protein Foundation Models

### v1

@@ Abstract
11footnotetext: Core contributor.22footnotetext: Corresponding authors.44footnotetext: Work completed while an exchange student at Princeton University.
@@ 1 Introduction
As a central component of the central dogma, proteins play essential roles in almost all biological processes, such as antibody-mediated antigen recognition and immune defense Tonegawa1983 ; KÖHLER1975 . Given the critical role of proteins, extensive research efforts in the past have focused on protein understanding and design doi:10.1126/science.181.4096.223 ; doi:10.1126/science.1219021 ; Huang2016 . This progress has been significantly accelerated by the emergence of deep learning. The AlphaFold alphafold2 ; alphafold3 and RoseTTAFold series rosettafold2021 ; rosettafold2024 provide high accuracy predictions of protein–molecule interaction structures through all-atom modeling. The ESM2 esm2 series offers sequence-level representations of proteins, advancing the understanding of protein semantics. In parallel, methods such as RFdiffusion series rfdiffusion2023 ; ahern2025rfdiffusion2 enable de novo protein design via diffusion-based generative modeling. Building on these advances, foundation models like ESM3 esm3 extend these capabilities by incorporating multimodal information, enabling protein prediction and generation under diverse constraints and guidance.
These advances deepen our understanding of proteins but also raise potential dual-use concerns, as they could be used to design proteins that pose biosecurity risks. It remains unclear whether protein foundation models are susceptible to be attacked. Growing concerns about potential misuse have drawn significant attention from the academic community. In response, leading scientists launched the Responsible AI $\times$ Biodesign initiative to promote protective measures [^1] https://responsiblebiodesign.ai/. However, a comprehensive framework to prevent the misuse of biological models remains absent, making its development a shared priority among researchers Wang2025 . This includes the need for effective governance over the dual-use capabilities of biological AI models bloomfield2024ai .
To help advance this progress, we conduct a preliminary study here to evaluate the potential misuse risks associated with previously proposed protein foundation models. Similar to the red-teaming evaluations used to test jailbreak attempts in large language models (LLMs) such as LLAMA lermen2024lorafinetuningefficientlyundoes , we adapt this adversarial testing paradigm for protein foundation models. Due to the difficulty of directly assessing the harmfulness of entirely novel proteins, we specialize the red-teaming evaluation into a masked recovery task. This allows us to test whether the protein foundation models can recover the complete masked sequence and structure under appropriate generation strategies. It also helps to evaluate whether the model is capable of understanding and generating the core functional domains of harmful proteins.
@@
Here, we propose SafeProtein, the first red-teaming framework designed for protein foundation models to the best of our knowledge. SafeProtein consists of two key components: (1) a systematic red-teaming methodology that integrates sequence- and structure-based input prompts, incorporates the non-pathogenic Foldseek structural similarity search strategy, and employs a score function guided heuristic beam search to enable comprehensive adversarial evaluation; and (2) a carefully constructed red-teaming benchmark, SafeProtein-Bench, which includes a manually curated dataset of viral and toxin proteins, multiple sequence-masking strategies, and an evaluation pipeline. The evaluation pipeline determines jailbreak success by jointly assessing sequence and structural similarity of the generated proteins. By conducting red-teaming exercises on protein foundation models, SafeProtein aims to uncover vulnerabilities and provide guidance for the development of stronger protective techniques.
We summarize our main contributions as follows:
@@
  • SafeProtein-Bench: the first dedicated benchmark for protein red-teaming, including a curated dataset and a comprehensive evaluation protocol.
@@ 2 SafeProtein: Red-Teaming Methodology for Protein Models
In this paper, the core objective of red-teaming a protein foundation model (Protein-FM) is to design a set of input prompts and generation schemes to test whether Protein-FM is capable of understanding and generating protein sequences or structures that are pathogenic, harmful, or otherwise biosecurity-relevant to living organisms.
Formally, consider a target Protein-FM and a judge function $\mathrm{JUDGE}$ that determines whether a generated protein corresponds to a harmful biological target in a database $\mathcal{D}$, based on sequence similarity, structural similarity, pathogen classification, or functional prediction. The red-teaming process can be formalized as:
@@ 2.2 Implementation Details of Generation Strategies
where $p_{\theta}(\cdot)$ is the parameterized reverse transition kernel of the diffusion model. For all positions where $m_{i}=1$, we impose the hard constraint $x_{i,t-1}=x^{\mathrm{cond}}_{i}$. The overall generation distribution is then given by:
@@ 3.1 Benchmark Dataset Construction
To evaluate the ability of models to reconstruct the structures of harmful proteins in red-teaming experiments, we retain only proteins with experimentally determined crystal structures. These proteins are generally well-studied, and their functional domains have been clearly characterized. Following common practice in protein foundation and generative models esm3 ; wang2025dplm ; geffner2025proteina , we further exclude entries shorter than 30 or longer than 1000 amino acids. Each entry in the dataset is accompanied by a detailed JSON file that records its sequence information, structural data, conservation profile, and the constructed masked-sequence inputs. The final curated dataset contains 429 proteins (Figure 1), and all entries were manually inspected. A complete list of toxin-producing organisms and regulated viruses included in the benchmark dataset is provided in Appendix A.
Figure 1: Details of benchmark dataset. (A) Taxonomic distribution of the test cases in our dataset. (B) Length distribution of the sequences in dataset.
@@ 4.5 Ablation Analysis, Limitations, and Future Work
Due to the limitation that larger-scale ESM3 models are only accessible via the closed and costly API, we were unable to test them directly. In addition, we adopted a more stringent joint sequence–structure criterion to reduce false positives, as we could not directly determine whether the generated proteins maintained similar functionality solely based on the structural similarity. Future work could involve laboratory validation of the generated sequences and collaboration with developers and communities working on advanced protein foundation models to conduct red-teaming tests on larger-scale models.
@@ 5 Conclusions and Ethics Statement
Appendix A SafeProtein-Bench Dataset Lists
Table S1: List of toxin-producing organisms in the benchmark.

### v2

@@ Abstract
22footnotetext: Corresponding authors.44footnotetext: Work was done when JF, $Z$ and RJ were visiting Princeton University.
Figure 1: Overview of (A) SafeProtein and (B) SafeProtein-Bench.
@@ 1 Introduction
As a central component of the central dogma, proteins play essential roles in almost all biological processes, such as antibody-mediated antigen recognition and immune defense Tonegawa1983 ; KÖHLER1975 . Given the critical role of proteins, extensive research efforts in the past have focused on protein understanding and design doi:10.1126/science.181.4096.223 ; doi:10.1126/science.1219021 ; Huang2016 . This progress has been significantly accelerated by the emergence of deep learning. The AlphaFold alphafold2 ; alphafold3 and RoseTTAFold series rosettafold2021 ; rosettafold2024 provide high accuracy predictions of protein–molecule interaction structures through all-atom modeling. The ESM2 esm2 series offers sequence-level representations of proteins, advancing the understanding of protein semantics. In parallel, methods such as RFdiffusion series rfdiffusion2023 ; ahern2025rfdiffusion2 ; butcher2025novo enable de novo protein design via diffusion-based generative modeling. Building on these advances, foundation models like ESM3 esm3 extend these capabilities by incorporating multimodal information, enabling protein prediction and generation under diverse constraints and guidance.
These advances deepen our understanding of proteins but also raise potential dual-use concerns, as they could be used to design proteins that pose biosecurity risks. This is because these foundation models may have already internalized the ability to understand and generate harmful proteins during training. It remains unclear whether protein foundation models are susceptible to be attacked. Growing concerns about potential misuse have drawn significant attention from the academic community. In response, leading scientists launched the Responsible AI $\times$ Biodesign initiative to promote protective measures [^1] https://responsiblebiodesign.ai/. However, a comprehensive framework to prevent the misuse of biological models remains absent, making its development a shared priority among researchers Wang2025 . This includes the need for effective governance over the dual-use capabilities of biological AI models bloomfield2024ai .
To advance this progress, we conduct a preliminary study here to evaluate the potential misuse risks of previously proposed protein foundation models. Similar to the red-teaming evaluations used to test jailbreak attempts in large language models (LLMs) lermen2024lorafinetuningefficientlyundoes ; zou2023universaltransferableadversarialattacks ; mazeika2024harmbenchstandardizedevaluationframework ; zeng2024airbench2024safetybenchmark ; zhou2025autoredteamerautonomousredteaming , we adapt this adversarial testing paradigm for protein foundation models. Due to the difficulty of directly assessing the harmfulness of entirely novel proteins, we specialize the red-teaming evaluation into a masked recovery task. This allows us to test whether the protein foundation models can recover the masked sequence and structure under appropriate generation strategies. It also helps to evaluate whether the model is capable of understanding and generating the core functional domains of harmful proteins.
@@
Here, we propose SafeProtein (Figure 1A), the first red-teaming framework designed for protein foundation models to the best of our knowledge. SafeProtein consists of two key components: (1) a systematic red-teaming methodology that integrates sequence- and structure-based input prompts, incorporates the non-pathogenic Foldseek structural similarity search strategy, and employs a score function guided heuristic beam search to enable comprehensive adversarial evaluation; and (2) a carefully constructed red-teaming benchmark, SafeProtein-Bench (Figure 1B), which includes a manually curated dataset of viral and toxin proteins, multiple sequence-masking strategies, and an evaluation pipeline. The evaluation pipeline determines jailbreak success by jointly assessing sequence and structural similarity of the generated proteins. By conducting red-teaming exercises on protein foundation models, SafeProtein aims to uncover vulnerabilities and provide guidance for the development of stronger protective techniques. We summarize our main contributions as follows:
@@
  • SafeProtein-Bench: the first benchmark for protein red-teaming, including a curated dataset of 429 experimentally resolved harmful proteins and a dual-criteria evaluation protocol.
@@ 2 SafeProtein: Red-Teaming Methodology for Protein Models
In this paper, the core objective of red-teaming a protein foundation model (Protein-FM) is to design a set of input prompts and generation schemes to test whether Protein-FM is capable of understanding and generating protein sequences or structures that are pathogenic, harmful, or otherwise biosecurity-relevant to living organisms. Formally, consider a target Protein-FM and a judge function $\mathrm{JUDGE}$ that determines whether a generated protein corresponds to a harmful biological target in a database $\mathcal{D}$, based on sequence similarity, structural similarity, pathogen classification, or functional prediction. The red-teaming process can be formalized as:
@@ 2.2 Implementation Details of Generation Strategies
where $p_{\theta}(\cdot)$ is the parameterized reverse transition kernel. For all positions where $m_{i}=1$, we impose the hard constraint $x_{i,t-1}=x^{\mathrm{cond}}_{i}$. The overall generation distribution is then given by:
@@ 3.1 Benchmark Dataset Construction
To evaluate the ability of models to reconstruct the structures of harmful proteins in red-teaming experiments, we retain only proteins with experimentally determined crystal structures. These proteins are generally well-studied, and their functional domains have been clearly characterized. Following common practice in protein foundation and generative models esm3 ; wang2025dplm ; geffner2025proteina , we further exclude entries shorter than 30 or longer than 1000 amino acids. Each entry in the dataset is accompanied by a detailed JSON file that records its sequence information, structural data, conservation profile, and the constructed masked-sequence inputs. The final curated dataset contains 429 proteins (Figure S1), and all entries were manually inspected. A complete list of toxin-producing organisms and regulated viruses included in the benchmark dataset is provided in Tables S1 and S2.
@@ 4.5 Ablation Analysis, Limitations, and Future Work
Due to the limitation that larger-scale ESM3 models are only accessible via the closed and costly API, we were unable to test them directly. In addition, we adopted a more stringent joint sequence–structure criterion to reduce false positives, as we could not directly determine whether the generated proteins maintained similar functionality solely based on the structural similarity. Future work could involve laboratory validation of the generated sequences, the integration of jailbreak alignment techniques (e.g., KPO wang2025enhancing ) into model development, and collaboration with developers and communities working on advanced protein foundation models to conduct red-teaming tests on larger-scale models.
@@ 5 Conclusions and Ethics Statement
Appendix A SafeProtein-Bench Dataset Details
Table S1: Details of toxin-producing organisms in the benchmark.
@@ Appendix A SafeProtein-Bench Dataset Details
Figure S1: Details of SafeProtein-Bench dataset. (A) Taxonomic distribution of the test cases in our dataset. (B) Length distribution of the sequences in dataset.